OpenStudy (stamp):

For any integer n > 0, the triangular number Tn is the number of dots in a triangular array with n points on each side. What is the value of the 64th triangle number?

OpenStudy (stamp):

\[T_1=1, \ T_2=3,\ T_3=6, \ T_4=10\]\[Find \ T_{64}.\]

OpenStudy (anonymous):

Note that T(1)=1, T(2)+1+2, T(3)=1+2+3... So we have T(n)=1+2+3+4+...+(n-1)+n

OpenStudy (stamp):

@Xavier Correct. Where do we go from here?

OpenStudy (anonymous):

This is generally a well known sum that evaluates to n(n+1)/2 The way this can be shown is that you can make n/2 (n+1)s by pairing the last and the first number, then the second and the second last number etc. So: n+1 2+(n-1) 3+(n-3) ... We can do this n/2 times so the sum evaluates to n(n+1)/2 Or it can be proven inductively I guess

OpenStudy (anonymous):

Sorry the second one in that small list should be 1+(n-1) then 2+(n-2)

OpenStudy (stamp):

@Xavier You are correct. I was considering this problem earlier and could not come up with a solution. I referenced some resources and ended up finding an approach, but it was not as ideal as yours. Thank you for your input.

OpenStudy (anonymous):

Glad to help.

OpenStudy (whpalmer4):

Here's another way of looking at this: Your triangle with \(n\) dots on a side is also \(n\) layers tall. If you take another such triangle (same value of \(n\)), flip it upside down, and nestle the two together side-by-side, you have a rectangular array of dots which is \(n*(n+1)\) dots in total. The triangular number is half of those dots, so \[T_n = \frac{n(n+1)}{2}\] @Xavier's explanation is the same trick that Gauss used to quickly sum the numbers from 1 to 100 as a child. Read a hundred or so versions of the story at http://www.sigmaxi.org/amscionline/gauss-snippets.html